A piece of wire is shaped to enclose a square with an area of 121 in^2. It is reshaped to enclose a rectangle which has a length of 13 in. What is the area of the triangle?
Answer by shweta(56)

Wire is shaped to enclose a square with area 121 inch^2
Let each side of the square be x inches
Area of square, A = x^2 = 121
x=sqrt(121)
x=11
Length of the wire will be same as the boundary of the square
Therefore,
Length of the wire= Perimeter of the square
= 4* 11
=44 inches
Now same wire is reshaped to enclose a rectangle of length 13 inches,so length of the wire will be same as the perimeter of the rectangle
Therefore, perimeter of rectangle = 44
Let 'l' be the length and 'b' be breadth of the rectangle
Perimeter of rectangle= 2*(l+b)=44
2*(13+b)=44
13+b=44/2
13+b=22
b=22-13
b=9 inch
Area of the rectangle= l*b
13*9= 117 inch^2